OpenStudy (anonymous):

The number of bacteria present in a culture after t minutes is given as b=100e^kt There are 6639 bacteria present after 5 minutes. Find k. Explain how you solve this problem.

OpenStudy (anonymous):

i guess you are supposed to solve \[\large 6639=100e^{5k}\]for \(k\)

OpenStudy (anonymous):

the explanation is that you know what \(b\) is when \(t=5\)

OpenStudy (anonymous):

do you know how to solve it? it take only 3 steps

OpenStudy (anonymous):

no :(

OpenStudy (anonymous):

ok lets to it step by step in case you have to do it on your own some time

OpenStudy (anonymous):

step 1 is divide by 100

OpenStudy (anonymous):

you get \[\large 66.39=e^{5t}\] good so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so log66.39=5t?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

t=.37?

OpenStudy (anonymous):

no k is .37

OpenStudy (anonymous):

i didn't check it is \[\frac{\ln(66.39)}{5}\] let me check it

OpenStudy (anonymous):

that is not what i get i take it you are using a calculator

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what kind? if you don't have to do this on a test i can show you here, but if you do we should make sure you can do it correctly on your calculator

OpenStudy (anonymous):

just google calc

OpenStudy (anonymous):

i get .36 actually

OpenStudy (anonymous):

...weird

OpenStudy (anonymous):

send me the link

OpenStudy (anonymous):

https://www.google.com/#q=ln%2866.39%29%2F5

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i am going to guess you did not end the parentheses just a guess

OpenStudy (anonymous):

idk...what happened there

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

thought so

OpenStudy (anonymous):

can you help me with another?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

The time required to grow a certain bacteria in a culture beginning with 100 bacteria is lnb-ln100/1.532, where B is the number of bacteria and t is the time in hours. How much time is required to grow a culture of 15,600 bacteria? Show the steps you used to find the number of hours.

OpenStudy (anonymous):

there is no t in your expression

OpenStudy (anonymous):

t=lnb-ln100/1.532

OpenStudy (anonymous):

oh ok so i guess we have to solve \[t=\frac{\ln(1500)-\ln(100)}{1.523}\] if i am reading your expression correctly that is just a straight up calculation nothing really so solve for

OpenStudy (anonymous):

it is the whole thing over \(1.523\) right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

look reasonable?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

k that was even easier, a direct substitution

OpenStudy (anonymous):

thank you sooo much

OpenStudy (anonymous):

youve been very helpful

OpenStudy (anonymous):

your quite welcome